I need to answer using addition, subtraction or substitution for the following problem:
a - 2b = 2
a + b = 2(b+3)

Answer by praseenakos@yahoo.com(507)

QUESTION:

a - 2b = 2
a + b = 2(b+3)

ANSWER:

a - 2b = 2 -----------(1)

a + b = 2(b+3)

==> a + b = 2b + 6

==> a + b - 2b = 6

==> a - b = 6 ----(2)

(1) - (2) ==>
==> a - 2b - (a- b) = 2 - 6

==> a - 2b - a + b = -4

==> -b = -4

==> b = 4

Now substitute the value b = 4 in (2)

==> a - 4 = 6

Add 4 on both sides,

==> a = 6 + 4

==> a = 10

So the solution is a = 10 and b = 4


Hope you understood.


Regards.

Praseena.